Maximize:4x1+12x2

subject to:

3x1+x2<=180

x1+2x2<=100

-2x1+2x2<=40

X1>=0

x2>=0

Find the optimal solution for the above model


1
Expert's answer
2021-01-19T01:54:32-0500

To find the optimal solution we have to draw the graph of the graph of the given constraints making all the given inequation as equation such as :

"L_1:3x_1+x_2=180"

"L_2:x_1+2_2=100"

"L_3:-2x_1+2x_2=40"

Now we draw the graph as follows:




Feasible solution of the given L.P.P was shown in the shaded region "OABCDO" .

Since the shaded region is bounded, each corner points gives the optimal solution of the given L.P.P.

Solving "L_1" and "L_2" we get the corner point "B(52,24)"

Solving "L_1" and "L_3" we get the corner point "C(20,40)"

And also "A(60,0)" , "D(0,20)" and "O(0,0)" are the corner points of the feasible region.

Now at these corner points we will find the value of "Z=4x_1+12x_2"

Hence "Z_{(60,0)}=(4\u00d760)+(12\u00d70)=240"

"Z_{(54,24)}=(4\u00d754)+(12\u00d724)=504"

"Z_{(20,40)}=(4\u00d720)+(12\u00d740)=560"

"Z_{(0,20)}=(4\u00d70)+(12\u00d720)=240"

Here we see that at "C(20,40)" , "Z=4x_1+12x_2" gives the maximum value.

Therefore "x_1=20" and "x_2=40" is the optimal solution of the above model.